### v1

# Unsupervised Evolutionary Cell Type Matching via Entropy-Minimized Optimal Transport

Abstract
Identifying evolutionary correspondences between cell types across species is a fundamental challenge in comparative genomics and evolutionary biology. Existing approaches often rely on either reference-based matching, which imposes asymmetry by designating one species as the reference, or projection-based matching, which may increase computational complexity and obscure biological interpretability at the cell-type level. Here, we present OT-MESH, an unsupervised computational framework leveraging entropy-regularized optimal transport (OT) to systematically determine cross-species cell type homologies. Our method uniquely integrates the Minimize Entropy of Sinkhorn (MESH) technique to refine the OT plan. It begins by selecting genes with high Signal-to-Noise Ratio (SNR) to capture the most informative features, from which a cost matrix is constructed using cosine distances between cell-type centroids. Importantly, the MESH procedure iteratively refines the cost matrix, leading to a transport plan with significantly enhanced sparsity and interpretability of the resulting correspondence matrices. Applied to retinal bipolar cells (BCs) and retinal ganglion cells (RGCs) from mouse and macaque, OT-MESH accurately recovers known evolutionary relationships and uncovers novel correspondences, one of which was independently validated experimentally. Thus, our framework offers a principled, scalable, symmetric, and interpretable solution for evolutionary cell type mapping, facilitating deeper insights into cellular specialization and conservation across species.
1 Introduction
Understanding evolutionary relationships between cell types across species is central to elucidating mechanisms underlying cellular specialization and diversity. Recent advances in single-cell RNA sequencing (scRNA-seq) have greatly improved the characterization of cell types based on their gene expression profiles across diverse species wangTracingCelltypeEvolution2021 ; hahnEvolutionNeuronalCell2023a . However, a fundamental challenge persists: given independently identified clusters of transcriptionally similar cells from different species, how can we systematically and reliably determine their evolutionary correspondences?
Existing methods for cross-species cell type alignment broadly fall into two categories shaferCrossSpeciesAnalysisSingleCell2019 : (1) reference-based matching, in which classifiers are trained on labeled data from a reference species to predict cell types in a query species shekharCOMPREHENSIVECLASSIFICATIONRETINAL2016a ; pengMolecularClassificationComparative2019a ; zhangEvolutionaryDevelopmentalSpecialization2024 ; wangMolecularCharacterizationSea2024 ; and (2) projection-based matching, in which individual-cell gene expression profiles from different species are embedded into a shared low-dimensional space using algorithms such as Canonical Correlation Analysis (CCA), Harmony, or other manifold alignment techniques butlerIntegratingSinglecellTranscriptomic2018b ; korsunskyFastSensitiveAccurate2019 ; rosenUniversalCellEmbeddings2024a . Reference-based matching inherently introduces asymmetry by designating one species as the "reference" and the other as the "query," potentially biasing outcomes, especially when species differ significantly in cell type numbers or evolutionary distances. Projection-based matching, though symmetric, typically operates at the level of individual cells, resulting in high computational demands and potentially obscuring direct biological interpretability of cell-type level relationships. Alternatively, cell type correspondences have been inferred through expert-curated morphological, physiological, or marker-gene criteria hahnEvolutionNeuronalCell2023a ; badenVertebrateRetinaWindow2024 . While valuable, these expert-driven methods depend heavily on prior hypotheses and manual validations, limiting scalability and systematic applicability.
To address these limitations, we introduce OT-MESH, an unsupervised computational framework. It combines entropy-regularized optimal transport (OT), a powerful tool for comparing distributions and finding correspondences, with a novel integration of the Minimize Entropy of Sinkhorn (MESH) refinement strategy zhangUnlockingSlotAttention2023 . OT is well-suited for this task due to its inherent ability to handle asymmetric numbers of types and partial correspondences. However, standard entropy-regularized OT often yields diffuse, non-sparse transport plans, hindering direct biological interpretation. Inspired by developments in machine learning and computer vision locatelloObjectCentricLearningSlot2020 ; sanderSinkformersTransformersDoubly2022 , we adapt the MESH technique to iteratively refine the OT cost matrix, promoting sparsity in the cross-species cell type alignment. Our method begins with Signal-to-Noise Ratio (SNR) based gene selection to identify informative features. It then constructs a cost matrix from cosine distances between cell-type centroids in this selected gene space. The MESH procedure is subsequently applied to refine the transport plan, yielding sparse and interpretable correspondence matrices.
We illustrate our approach using established datasets of bipolar cells (BCs) and retinal ganglion cells (RGCs) from mouse and macaque retinas hahnEvolutionNeuronalCell2023a ; pengMolecularClassificationComparative2019a ; tranSinglecellProfilesRetinal2019a . Leveraging existing knowledge for validation, we demonstrate that our method consistently identifies known evolutionary correspondences and discovers novel cell-type homologies, including one correspondence independently validated experimentally wangONtypeDirectionselectiveGanglion2023 . Thus, OT-MESH provides a principled, scalable, and interpretable computational framework, enhancing our ability to map cell type evolution and advance its understanding.
2 Background
2.1 Optimal Transport (OT) for Correspondence Problems
OT provides a principled mathematical framework for identifying correspondence between distributions. Given two distributions supported on spaces $\mathbf{X}$ and $\mathbf{Y}$, OT seeks the optimal mapping that minimizes the total cost of transporting mass from $\mathbf{X}$ to $\mathbf{Y}$. In discrete form, it involves finding a transport plan $\mathbf{W}$ that minimizes
$L(\mathbf{W})=\sum_{ij}W_{ij}C_{ij},$ (1)
where $C_{ij}$ is the cost of transporting unit mass from point $i$ in $\mathbf{X}$ to point $j$ in $\mathbf{Y}$.
To enhance computational tractability and numerical stability, OT problems often incorporate entropy regularization, which encourages smoother transport plans. The entropy-regularized OT problem is formulated as:
$L(\mathbf{W})=\sum_{ij}W_{ij}C_{ij}+\alpha\sum_{ij}W_{ij}\Bigl{(}\log W_{ij}-1\Bigr{)},$ (2)
subject to the marginal constraints:
$\sum_{j}W_{ij}=p_{i},\quad\sum_{i}W_{ij}=q_{j}.$ (3)
Here, $p_{i}$ and $q_{j}$ define prescribed marginals. This entropy-regularized problem can be efficiently solved using the Sinkhorn algorithm cuturiSinkhornDistancesLightspeed2013 .
2.1.1 Sinkhorn Algorithm
Given a cost matrix $\mathbf{C}\in\mathbb{R}^{a\times b}$, a regularization parameter $\alpha>0$, and prescribed marginal distributions $\mathbf{p}\in\mathbb{R}^{a}$ and $\mathbf{q}\in\mathbb{R}^{b}$, the Sinkhorn algorithm computes the Gibbs kernel as:
$\mathbf{G}=\exp\left(-\frac{\mathbf{C}}{\alpha}\right).$ (4)
It iteratively updates two scaling vectors $\mathbf{u}$ and $\mathbf{v}$ as follows:
$\mathbf{u}\leftarrow\mathbf{p}\oslash(\mathbf{G}\mathbf{v}),\qquad\mathbf{v}\leftarrow\mathbf{q}\oslash(\mathbf{G}^{\top}\mathbf{u}),$ (5)
[…]
$\mathbf{W}=\operatorname{diag}(\mathbf{u})\,\mathbf{G}\,\operatorname{diag}(\mathbf{v}).$ (6)
In practice, computations are frequently performed in log-space to maintain numerical stability.
2.2 Minimize Entropy of Sinkhorn (MESH)
While entropy regularization makes OT computationally tractable, it often produces overly diffuse (low interpretability) transport plans with limited interpretability, especially when the underlying biological correspondences are expected to be sparse. To address this, the MESH approach zhangUnlockingSlotAttention2023 iteratively refines the cost matrix itself to minimize the entropy of the resulting transport plan:
$\mathbf{C}^{\prime}(t+1)=\mathbf{C}^{\prime}(t)-\lambda\frac{\nabla_{\mathbf{C}^{\prime}(t)}H\bigl{(}\text{sinkhorn}(\mathbf{C}^{\prime}(t))\bigr{)}}{\|\nabla_{\mathbf{C}^{\prime}(t)}H\bigl{(}\text{sinkhorn}(\mathbf{C}^{\prime}(t))\bigr{)}\|},$ (7)
where the entropy of the transport plan $\mathbf{W}(t)=\text{sinkhorn}(\mathbf{C}^{\prime}(t))$ is defined as:
$H(\mathbf{W}(t))=-\sum_{ij}W_{ij}(t)\log W_{ij}(t),$ (8)
and $\lambda$ is a learning rate. This iterative approach yields sparse and interpretable correspondence matrices by concentrating mass onto fewer, more meaningful connections.
2.3 OT Applications in Single-Cell Biology
OT methods have found increasing applications in single-cell biology, particularly for aligning cell states across experimental conditions, measurement modalities, or developmental time points demetciSCOTSingleCellMultiOmics2022 ; gossiMatchingSingleCells2023 ; penaConstructingCelltypeTaxonomy2025 . A key advantage of OT approaches lies in their natural ability to handle partial or incomplete correspondences. Recent studies have leveraged OT to construct hierarchical cell-type taxonomies, effectively identifying cell types uniquely represented in specific datasets penaConstructingCelltypeTaxonomy2025 . Nonetheless, OT-based methods have yet to be comprehensively developed and validated specifically for cross-species evolutionary cell type comparisons. In this study, we address this critical gap by adapting and extending OT methodologies, particularly through the MESH procedure, to the problem of cross-species cell type matching.
3 Methods
3.1 Problem Formulation
We consider two species containing $a$ and $b$ distinct cell types, indexed by $i$ and $j$ respectively. For the $i$th cell type in the first species, let $\mathbf{x}_{ik}\in\mathbb{R}^{1\times g}$ represent the gene expression vector of the the $k$th cell; similarly, let $\mathbf{y}_{jl}\in\mathbb{R}^{1\times g}$ denote expression for the $l$th cell of the $j$th type in the second species. Our goal is to identify an optimal correspondence matrix $\mathbf{W}\in\mathbb{R}^{a\times b}$ where each element $W_{ij}$ assigns a probability to each pair of cell types being evolutionarily related.
3.2 Gene Selection by Signal-to-Noise Ratio (SNR)
Rather than using all available genes, which can introduce noise and increase computational burden, we identify a subset of highly informative genes based on their expression SNR hahnEvolutionNeuronalCell2023a . SNR prioritizes genes that show high variance between cell types relative to variance within cell types, thus highlighting genes that effectively distinguish types. Specifically, for each gene $g$, intra-type variances $s^{2}_{i,g}$ and $s^{2}_{j,g}$ are computed as:
$s^{2}_{i,g}=\frac{\sum_{k}\left(x_{ik,g}-\bar{x}_{i:,g}\right)^{2}}{n_{i}},\quad s^{2}_{j,g}=\frac{\sum_{l}\left(y_{jl,g}-\bar{y}_{j:,g}\right)^{2}}{n_{j}}.$ (9)
where $\bar{x}_{i:,g}$ and $\bar{y}_{j:,g}$ denote the mean expression level of gene $g$ for cell types $i$ and $j$ respectively, and $n_{i}$, $n_{j}$ represent the numbers of cells in those types. The SNR for each gene $g$ is calculated by:
$\text{SNR}(g)=\frac{\mathrm{Var}\left(\{\bar{x}_{1:,g},\dots,\bar{x}_{a:,g},\bar{y}_{1:,g},\dots,\bar{y}_{b:,g}\}\right)}{\frac{1}{a+b}\left(\sum_{i=1}^{a}{s^{2}_{i,g}}+\sum_{j=1}^{b}{s^{2}_{j,g}}\right)}.$ (10)
We retain genes exceeding a predefined SNR threshold, resulting in $m$ selected genes used to construct the centroid matrices $\hat{\mathbf{X}}\in\mathbb{R}^{a\times m}$ for the first species, and $\hat{\mathbf{Y}}\in\mathbb{R}^{b\times m}$ for the second species.
3.3 Cost Matrix Construction
We define a scale-invariant cost matrix $\mathbf{C}\in\mathbb{R}^{a\times b}$ based on the cosine distance between cell-type centroids:
$C_{ij}=1-\frac{\hat{\mathbf{X}}_{i}\cdot\hat{\mathbf{Y}}_{j}}{\|\hat{\mathbf{X}}_{i}\|\,\|\hat{\mathbf{Y}}_{j}\|}.$ (11)
Here, $\hat{\mathbf{X}}_{i}$ and $\hat{\mathbf{Y}}_{j}$ are the centroid vectors of cell types $i$ and $j$ respectively.
3.4 Entropy-Regularized OT with MESH
To obtain sparse and interpretable correspondences, we employ entropy-regularized OT refined by the MESH procedure. We begin with a slightly perturbed initial cost matrix zhangUnlockingSlotAttention2023 :
$\mathbf{C}^{\prime}(0)=\mathbf{C}+\varepsilon,\quad\varepsilon_{ij}\sim\mathcal{N}(0,10^{-6}),$ (12)
At each iteration $t$, we compute the transport plan $\mathbf{W}(t)$ from the current cost matrix $\mathbf{C}^{\prime}(t)$ using the Sinkhorn algorithm and then quantify its entropy:
$F\bigl{(}\mathbf{C}^{\prime}(t)\bigr{)}=H(\mathbf{W}(t))=-\sum_{ij}W_{ij}(t)\log W_{ij}(t).$ (13)
The cost matrix is updated by gradient descent, normalized to maintain stable updates:
$\mathbf{C}^{\prime}(t+1)=\mathbf{C}^{\prime}(t)-\lambda\frac{\nabla_{\mathbf{C}^{\prime}(t)}F\bigl{(}\mathbf{C}^{\prime}(t)\bigr{)}}{\|\nabla_{\mathbf{C}^{\prime}(t)}F\bigl{(}\mathbf{C}^{\prime}(t)\bigr{)}\|},$ (14)
After $T$ iterations, the resulting cost matrix $\mathbf{C}^{\prime}(T)$ is used to compute the final transport plan:
$\mathbf{W}^{*}=\text{sinkhorn}(\mathbf{C}^{\prime}(T)).$ (15)
3.5 Algorithm Summary
Our entire approach is summarized as Algorithm 1 and illustrated schematically in Figure 1.
Algorithm 1 Evolutionary Cell Type Matching via Entropy-Regularized OT with MESH
1:Input: Centroid matrices from SNR-selected genes, $\hat{\mathbf{X}}\in\mathbb{R}^{a\times m}$ and $\hat{\mathbf{Y}}\in\mathbb{R}^{b\times m}$, regularization parameter $\alpha$, MESH learning rate $\lambda$, number of MESH iterations $T$
2:Compute cost matrix $\mathbf{C}\in\mathbb{R}^{a\times b}$ as in Eq. 11
3:Initialize marginal distributions $\mathbf{p}$, $\mathbf{q}$, and $\mathbf{C}^{\prime}(0)$ as in Eq. 12
4:For $t=0$ to $T-1$:
5:      Compute $\mathbf{W}(t)=\text{sinkhorn}(\mathbf{C}^{\prime}(t))$
6:      Calculate entropy $F\bigl{(}\mathbf{C}^{\prime}(t)\bigr{)}$ as in Eq. 13
7:      Update $\mathbf{C}^{\prime}(t+1)$ using Eq. 14
8:Output: Final correspondence matrix $\mathbf{W}^{*}$ as in Eq. 15
Figure 1: Illustration of the proposed method. Cell type centroids are computed using SNR-selected genes, and a cost matrix is constructed using cosine distance between these centroids for the first and second species. The MESH procedure iteratively refines this cost matrix to produce the final correspondence matrix that reveals evolutionarily related cell types with sparse, interpretable mappings that quantify the strength of relationships between cell types across species.
4 Experimental Setup
4.1 Datasets
We evaluated our method using three established retinal cell datasets:
We first validated our method on macaque retinal BC types with known correspondences between peripheral and foveal regions. Specifically, we created a ground-truth test case comprising 12 BC types with exact 1:1 matching across these regions (excluding the fovea-specific "OFFx" type). Subsequently, we extended our analyses to cross-species comparisons: mouse BC (15 types) versus macaque BC (12 types), and mouse RGC (45 types) versus macaque foveal RGC (18 types).
4.2 Preprocessing and Feature Selection
We adhered to previously established preprocessing and feature selection protocols hahnEvolutionNeuronalCell2023a ; shekharCOMPREHENSIVECLASSIFICATIONRETINAL2016a ; pengMolecularClassificationComparative2019a ; tranSinglecellProfilesRetinal2019a ; qiaoFactorizedDiscriminantAnalysis2023c ; qiaoDecipheringGeneticCode2024a . Orthologous genes between mouse and macaque were identified using standard orthology databases, resulting in 10,416 common genes. Cell transcript counts were normalized to the median total transcript count per cell, followed by log-transformation to stabilize variance and mitigate technical variability.
For feature selection, we ranked genes by their SNR (Eq. 10) and retained genes surpassing an empirical threshold (SNR > 100, yielding a manageable number of highly discriminative genes, typically a few hundred to a thousand in our datasets), resulting in a subset of genes that effectively differentiate cell types. Cell-type centroids were computed based on these selected genes, forming input matrices $\hat{\mathbf{X}}$ and $\hat{\mathbf{Y}}$ to the OT algorithm.
4.3 Cell Type Correspondence
We formulated cell type correspondence as an OT problem with uniform marginals, $p_{i}=1/a$ and $q_{j}=1/b$. This choice reflects an agnostic prior about correspondences and prevents dominant cell types from disproportionately influencing the correspondence matrix.
To quantify correspondence strength between cell types between cell types $i,j$ from the final transport matrix $\mathbf{W}^{*}$, we defined an alignment score $s_{ij}$ as:
$s_{ij}=\frac{1}{2}\left(\frac{W^{*}_{ij}}{\sum_{k=1}^{a}{W^{*}_{kj}}}+\frac{W^{*}_{ij}}{\sum_{l=1}^{b}{W^{*}_{il}}}\right).$ (16)
This score $s_{ij}$ represents the average of two probabilities: the probability of cell type $i$ in species one matching type $j$ among all cell types in species two, and vice versa. By averaging both row-wise and column-wise normalized weights, this metric robustly identifies and ranks strong bidirectional evolutionary relationships.
For interpretation of cross-species RGC correspondences, we integrated established morphological and functional annotations from prior comprehensive studies qiaoDecipheringGeneticCode2024a ; goetzUnifiedClassificationMouse2022a . This allowed mapping of transcriptomic clusters from the mouse dataset tranSinglecellProfilesRetinal2019a to functionally characterized RGC subtypes, such as alpha, direction-selective, and intrinsically photosensitive RGCs, thereby facilitating biological interpretation of identified macaque RGC counterparts.
4.3.1 Evaluation Metrics
We used four key metrics to assess our approach:
  • Sparseness Score: Proportion of negligible entries (<0.001) in the correspondence matrix, indicating interpretability through sparsity.
  • Entropy: Quantifies uncertainty or diffuseness of correspondence assignments (Eq. 13); lower entropy indicates clearer correspondences.
  • Adjusted Rand Index (ARI): Measures agreement between predicted correspondences and known ground-truth assignments; higher ARI indicates greater accuracy.
  • Running Time: Computational efficiency measured by elapsed processing time.
4.3.2 Parameter Selection
We conducted systematic parameter optimization through grid search over the following parameter ranges:
  • Regularization parameter $\alpha$: [0.01, 0.1, 1.0, 10.0]
  • MESH learning rate $\lambda$: [0.1, 0.5, 1.0, 5.0]
  • Number of MESH iterations $T$: [1.0, 2.0, 4.0, 8.0]
For each parameter combination, we computed the transport loss (Eq. 2) and sparseness score. Because our primary aim was sparse and interpretable correspondences, we restricted our selection to parameter sets with sparseness scores above 0.5, choosing the set that minimized transport loss. Optimal parameters were independently determined for each experiment, detailed in SupplementaryFigures 1–3.
4.3.3 Comparative Baselines
To contextualize our method, we compared it against three baseline approaches:
  • XGBoost classification (reference-based matching): A supervised classification approach commonly used in previous cross-species studies pengMolecularClassificationComparative2019a ; zhangEvolutionaryDevelopmentalSpecialization2024 ; wangMolecularCharacterizationSea2024 . An XGBoost classifier is trained on cell types from a designated reference species using orthologous gene expression profiles, and applied to predict cell types in a query species. The correspondence matrix is derived by aggregating the predicted labels of individual query cells to quantify cell-type level relationships.
  • Harmony integration followed by 1-Nearest Neighbor (Harmony+1NN, projection-based matching): Harmony korsunskyFastSensitiveAccurate2019 integrates individual-cell transcriptomic profiles from different species into a shared low-dimensional embedding, correcting batch effects. Cell-type correspondences are inferred by assigning each cell to the cell type of its nearest neighbor from the other species in the integrated space. These individual-level assignments are then aggregated to form a correspondence matrix reflecting inter-species cell-type alignments.
  • Entropy-regularized OT without MESH refinement: To explicitly evaluate the benefit of MESH, we included standard entropy-regularized OT as a baseline, using the same SNR-selected genes for fair comparison.
All methods were evaluated using the previously described metrics, allowing rigorous comparison of accuracy, interpretability, and computational efficiency.
5 Results
5.1 Method Validation and Comparison
We first validated our OT-MESH method against known cell type correspondences between macaque peripheral and foveal bipolar cells (BCs), comparing its performance to three baseline methods: XGBoost, Harmony+1NN, and standard entropy-regularized OT without MESH. For XGBoost, the classifier was trained using macaque foveal BCs as the reference pengMolecularClassificationComparative2019a . For Harmony+1NN, we performed embedding based integration followed by cell type label transfer through nearest neighbor assignment in the harmonized PCA space korsunskyFastSensitiveAccurate2019 . Standard OT was performed using the same parameters as OT-MESH but without iterative entropy minimization.
Figure 2 illustrates the resulting correspondence matrices. The XGBoost approach (Figure 2A) yielded a predominantly diagonal structure with notable off-diagonal noise. Harmony+1NN (Figure 2B) improved upon XGBoost by producing cleaner correspondences, though some off-diagonal elements persisted. Standard OT without MESH refinement (Figure 2C) produced a highly diffuse correspondence matrix, making interpretation nearly impossible despite its computational speed. In contrast, OT-MESH (Figure 2D) resulted in a clean diagonal structure with high-confidence values, clearly recovering the true 1:1 correspondences.
Quantitative metrics (Table 1) reinforce these findings. OT-MESH achieved the best overall performance: highest sparseness (0.8678), lowest entropy (2.4597), and highest ARI (0.9779). Harmony+1NN was second in sparseness (0.8347) and ARI (0.9258) but was significantly slower (14.5306s vs. 0.7626s for OT-MESH). XGBoost had lower ARI (0.8871) and was slower than OT-MESH (2.3062s). Standard OT was fastest (0.0329s) but its extremely low sparseness (0), poor ARI (0.0240), and diffuse nature made it unsuitable for meaningful biological interpretation. These results demonstrate that OT-MESH provides a compelling balance of accuracy, interpretability, and computational efficiency for cross-species cell type correspondence.
Figure 2: Comparison of correspondence matrices for macaque peripheral to foveal BC type mapping. A) XGBoost shows diagonal structure but with some off-diagonal noise. B) Harmony+1NN improves alignment with cleaner structure. C) Regular OT produces a diffuse, uninterpretable mapping. D) OT-MESH yields the clearest diagonal structure with high-confidence 1:1 matchings, consistent with ground truth. Correspondence matrices are normalized so that the sum of all their entries is one.
Table 1: Performance comparison of cross-species cell type matching methods on macaque foveal vs. peripheral BC types. Best values for each metric are in bold.
Method Sparseness Score Entropy ARI Running Time (s)
XGBoost 0.8017 2.6738 0.8871 2.3062
Harmony+1NN 0.8347 2.5949 0.9258 14.5306
OT without MESH 0.0000 4.6934 0.0240 0.0329
OT-MESH 0.8678 2.4597 0.9779 0.7626
5.2 Cross-Species Correspondence of Bipolar Cell (BC) Types
We next applied OT-MESH to cross-species alignment of BC types between mouse (15 types) and macaque (12 types). The resulting correspondence matrix (Figure 3A) reveals clear evolutionary matches, forming strong diagonal-like patterns.
Alignment scores (SupplementaryTable 1) quantify these correspondences. The strongest match was between mouse RBC and macaque RB (rod BCs), with a score of 0.91. Among cone BCs, macaque DB1, DB2, DB3a, DB3b, DB6, IMB, and BB/GB* types were uniquely and strongly assigned to mouse BC2, BC4, BC3A, BC3B, BC6, BC7, and BC8/9 types, respectively, consistent with previous reports hahnEvolutionNeuronalCell2023a ; shekharCOMPREHENSIVECLASSIFICATIONRETINAL2016a .
The remaining three macaque BC types showed more complex mappings. Macaque FMB corresponded to both mouse BC1A and BC1B. Macaque DB4 mapped to both mouse BC5A and BC5D. Notably, macaque DB5* showed distributed correspondence with mouse BC5B and BC5C. The mapping of DB5* (a cone BC) to mouse BC5B/C (cone BCs) is more consistent with their shared functional (ON pathway) and morphological (axon terminal stratification in middle IPL) properties  eulerRetinalBipolarCells2014b ; tsukamotoBipolarCellsMacaque2016 compared with a previous report mapping it to mouse RBC (a rod BC) shekharCOMPREHENSIVECLASSIFICATIONRETINAL2016a . This highlights OT-MESH’s ability to uncover complex homologies.
5.3 Cross-Species Correspondence of Ganglion Cell (RGC) Types
When applied to the more complex case of RGC types (45 mouse types vs. 18 macaque foveal types), OT-MESH successfully recovered meaningful cross-species correspondences despite the substantial difference in type numbers and greater evolutionary divergence. Figure 3B visualizes these relationships, showing a sparser correspondence matrix compared to BCs, but with several strong matches between functionally similar types, reflecting known conserved pathways and also specialized lineages.
Most notably, our analysis confirmed the correspondence between mouse alpha RGCs and primate midget/parasol ganglion cells with high confidence (SupplementaryTable 2). Specifically, mouse OFF transient alpha (C45) maps strongly to macaque OFF parasol ganglion cells (PG-OFF) (score: 0.61), while ON sustained alpha (C43) corresponds to macaque ON midget ganglion cells (MG-ON) (score: 0.52). Similarly, mouse ON transient alpha (C41) and OFF sustained alpha (C42) map to macaque ON parasol (PG-ON) and OFF midget ganglion cells (MG-OFF) respectively (scores: 0.23 and 0.20). This provides additional support for the evolutionary relationship between these cell types, consistent with the recent report hahnEvolutionNeuronalCell2023a . The correspondence is particularly compelling as it respects both response polarity (ON vs. OFF) and kinetics (sustained vs. transient), suggesting conservation of fundamental visual processing channels across mammalian evolution.
Importantly, OT-MESH identified the strongest overall correspondence between macaque fRGC12 and mouse type C10 (an ON direction-selective RGC) with a high alignment score of 0.66 (SupplementaryTable 2). This finding suggests a previously underappreciated evolutionary relationship between these types. Notably, the existence of an ON direction-selective RGC in primate retina was independently confirmed by Wang et al. wangONtypeDirectionselectiveGanglion2023 , lending strong external validation to our method’s predictive power. We also detected robust correspondences between macaque fRGC14 and mouse melanopsin-expressing ipRGCs (C40 and C33, corresponding to M1 subtypes) (scores: 0.60 and 0.40), confirming the conservation of this non-image-forming visual pathway across mammals; and the mapping between macaque fRGC11 and mouse F-mini-ON cells (C3) (score: 0.22), suggesting potential evolutionary links between these functionally specialized types. The correspondence matrix also revealed that many mouse RGC types lacked clear macaque counterparts, suggesting evolutionary specialization in the murine lineage.
Collectively, these results illustrate that OT-MESH captures both well-known and novel cross-species relationships among RGCs. The increased complexity and divergence in RGC mappings, relative to BCs, align with expectations for output neurons under greater evolutionary selection pressure hahnEvolutionNeuronalCell2023a ; pengMolecularClassificationComparative2019a .
Figure 3: Cross-species correspondence matrices. A) Mouse-to-macaque BC correspondence showing clear mapping between evolutionarily conserved types. B) Mouse-to-macaque RGC correspondence revealing more complex relationships due to the difference in type numbers (45 mouse vs. 18 macaque types).
6 Discussion
We have presented OT-MESH, an unsupervised computational framework that systematically identifies evolutionary correspondences between cell types across species. The core innovation lies in the integration of the MESH procedure with an SNR-based gene selection and centroid-based OT cost matrix. This combination transforms standard OT outputs into sparse and biologically interpretable correspondence matrices. Applied to retinal neuron datasets from mouse and macaque, OT-MESH successfully recovered known homologies and revealed novel, biologically plausible correspondences. Its ability to quantify alignment strength and robustly handle differences in cell type numbers enhances its utility for cross-species comparative analysis. The observed patterns of conservation (e.g., in BCs) and divergence (e.g., more specialized RGCs in mouse) align with existing biological understanding, while newly predicted matchings offer clear avenues for future experimental validation.
Compared to existing methods, OT-MESH demonstrates several clear advantages. In contrast to reference-based classifiers such as XGBoost pengMolecularClassificationComparative2019a ; zhangEvolutionaryDevelopmentalSpecialization2024 ; wangMolecularCharacterizationSea2024 , OT-MESH is unsupervised and symmetric, requiring no designation of reference versus query species. By operating directly on cell-type centroids rather than individual cells, it achieves significant computational efficiency compared to cell-level projection-based matching methods such as Harmony korsunskyFastSensitiveAccurate2019 . Unlike standard entropy-regularized OT, which yields overly diffuse and uninterpretable transport plans, the MESH refinement produces sparse correspondence matrices that align closely with biological intuition and ground truth. This interpretability enables direct hypothesis generation, as exemplified by the high-confidence alignment between macaque fRGC12 and mouse ON direction-selective RGC (C10), which was independently validated by Wang et al. wangONtypeDirectionselectiveGanglion2023 . This finding illustrates how OT-MESH not only recovers known relationships but also predicts new ones with potential biological relevance.
While this study focused on transcriptomic data, our framework is readily extensible. Future versions could incorporate additional modalities, such as epigenomic states, spatial localization, or electrophysiological properties, to more robustly define cell types and resolve ambiguities that transcriptomics alone may not disentangle. Similarly, although SNR-based gene selection and cosine distance proved effective and computationally tractable, alternative feature selection strategies or dissimilarity metrics beyond SNR and cosine distance could also be explored. Beyond the retina, OT-MESH can be applied to explore evolutionary mapping of cell types across other organ systems (e.g., cortical neurons, immune cell lineages) and more distant species.
In conclusion, by enabling unsupervised, interpretable, and biologically grounded cell-type correspondence, OT-MESH can advance the field’s ability to trace the evolution of cellular diversity and specialization across species, offering a valuable tool for comparative genomics and evolutionary biology.
Supplementary Table 1: Top-ranked correspondences between macaque and mouse BC types quantified by alignment scores (Eq.16).
Macaque BC Types (Peng et al. 2019) Mouse BC Types (Shekhar et al. 2016) Alignment Score
RB RBC 0.9101
DB3a BC3A 0.8930
DB1 BC2 0.8731
DB2 BC4 0.8458
IMB BC7 0.8277
DB3b BC3B 0.8268
DB6 BC6 0.8172
BB/GB* BC8/9 0.7330
DB4 BC5D 0.6630
DB5* BC5B 0.6496
FMB BC1A 0.6261
DB5* BC5C 0.4649
DB4 BC5A 0.4108
FMB BC1B 0.4015
Supplementary Table 2: Top 10 correspondences between macaque and mouse RGC types quantified by alignment scores. Mouse RGC annotations from Tran et al. [15] and Goetz et al. [23].
Macaque RGC (Peng et al. 2019) Mouse RGC (Tran et al. 2019) Mouse RGC Annotation Alignment Score
fRGC12 C10 ON direction-selective 0.6616
PG_OFF C45 OFF transient alpha 0.6060
fRGC14 C40 M1 ipRGC 0.5967
MG_ON C43 ON sustained alpha 0.5242
fRGC5 C8 0.4058
fRGC14 C33 M1 ipRGC 0.3971
PG_OFF C34 0.3524
PG_ON C41 ON transient alpha 0.2345
fRGC11 C3 F-mini-ON 0.2156
MG_OFF C42 OFF sustained alpha 0.1995